OpenStudy (anonymous):

insert two arithmetic means between each pair. the pair is -27 and 33

OpenStudy (anonymous):

Let the two numbers be a and b your series: -27, a, b, 33 T4 - T1 = 33 - (-27) = 60 60/3 = 20 so d (common difference) = 20 so a = -27 + d and b = a + d a = -7 b = 13 Your new series: -27, -7, 13, 33